my math didn't explain this to me:
A line that passes through the origin and is parallel to x + y = 6.
thank you,
Destiny

Answer by EMStelley(208)

First of all, we need to remember that parallel lines have the same slope. So we need to determine the slope of the given line, x + y = 6.
If we rearrange this to put it in y= form, we have
y = -x + 6
So the slope is -1. Now, remember that to determine the equation of a line, we need a point and a slope. We have the slope, and we are told that the line must pass through the origin, which is (0,0), so we have our point also. Using point-slope form of a line, we have:
y - 0 = -1(x - 0)
y = -x
I hope that helps!